# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ 4 Accretion disc model
In this paper, the compact object is surrounded by an infinitely geometrically thin accretion disc orbiting in the equatorial plane taken to be at $\theta=\pi/2$ without any loss of generality. Our disc extends from an inner radius, $r_{\mathrm{inner}}$, given by the radial position of the event horizon, to a sufficiently large outer radius $r_{\mathrm{outer}}$ whose contribution to the image formation is negligible for our choice of quickly decreasing emission profiles.
@@ 4.1 Motion of the emitting material
We suppose that the emitters of mass $m_{\mathrm{e}}>0$ can have three possible types of dynamics: either the particles in the disc rotate around the black hole with a Keplerian velocity, they fall radially towards the central singularity or they possess a more realistic mixed velocity. A sketch of the two extreme velocity fields is reproduced in Figure 3. We define the specific energy $\varepsilon$ and specific angular momentum $\ell$ as
@@ 4.1.1 Keplerian velocity
If the disc is Keplerian, the accreted matter follows a circular geodesic motion above the innermost stable circular orbit (ISCO), below which we assume that it spirals down along timelike geodesics with constants of motion $\varepsilon$ and $\ell$ equal to those of the ISCO (Cunningham 1975). A general timelike 4-velocity $u^{\,\mu}=p^{\,\mu}/m_{\mathrm{e}}=(u^{t},u^{r},0,u^{\varphi})$ can be expressed with the conserved quantities $\varepsilon$ and $\ell$ and via the normalisation $u^{\,\mu}u_{\mu}=-1$. Noting by primes the derivatives with respect to $r$, the expressions of the specific energy and z-component of the angular momentum for the RZ metric are then (Cárdenas-Avendaño et al. 2019)
@@
The last stable orbit being marginally stable, $V^{\prime\prime}_{\mathrm{eff}}$, the second derivative of the effective potential, must vanish at the radius of the ISCO, $r_{\mathrm{ms}}$, which is then the real root of the following equation with no analytical solution (Rezzolla & Zhidenko 2014):
@@ 4.1.2 Radial infall
When the accreted matter is in radial free-fall towards the central compact object, the particle evolves in a constant-$\theta$ plane with vanishing specific angular momentum $\ell$. The conservation of the specific energy $\varepsilon$ and the normalisation of the 4-velocity lead to (Kocherlakota & Rezzolla 2022):
@@
We also neglect the radial velocity at infinity, so that:
@@ 4.1.3 Mixed velocity
Realistic RIAF discs are neither totally radially infalling nor Keplerian, but, according to more sophisticated GRMHD simulations (Narayan et al. 2012), they likely consist of a mixed velocity comprising a radial component and a sub-Keplerian part (Yuan et al. 2022; Begelman et al. 2022). Moreover, the sub-Keplerian motion is expected to be predominant for the disc surrounding M87* (Chatterjee et al. 2023).
To define the sub-Keplerian velocity $u_{\mathrm{sub}}$ we follow the same reasoning of Paragraph 4.1.1, but we multiply the quantity $\ell/\varepsilon$ by the ‘sub-Keplerianity’ factor $\xi\in(0,1]$, so that above the ISCO
$\varepsilon_{\mathrm{sub}}=\sqrt{\frac{N^{3}(r)}{N(r)-\xi^{2}rN^{\prime}(r)}}\>,\quad\ell_{\mathrm{sub}}=\xi\sqrt{\frac{r^{3}N^{\prime}(r)}{N(r)-\xi^{2}rN^{\prime}(r)}}$
[…]
@@
Finally, $u^{t}_{\mathrm{mix}}$ is found by unit-normalisation and $u^{\varphi}_{\mathrm{mix}}=\Omega_{\mathrm{mix}}u^{t}_{\mathrm{mix}}$.
We note that this realistic mixed velocity does not correspond to a geodesic motion contrary to the purely Keplerian and radial ones, that are retrieved, respectively, by imposing $\omega_{r}=\omega_{\varphi}=1$ with $\xi=1$ and $\omega_{r}=\omega_{\varphi}=0$. In other words, the mixed velocity prescription can account for non-gravitational forces, that are expected to affect the emitters’ trajectory.
@@ 4.2 Synchrotron radiation
We assume the disc to be optically and geometrically thin, which means considering the optical depth to be idealistically null. The first hypothesis is only an approximation for the present EHT images of M87* and particularly for the photon rings observations, for which the opacity effects may play a role (Vincent et al. 2022). The assumption holds much better at 345 GHz, as absorption effects decrease with increasing frequency, and so it is reasonable for the observations at higher frequency of the ngEHT (Doeleman et al. 2023) or BHEX (Johnson et al. 2024). In the absence of absorption, the quantity $I_{\nu}/\nu^{3}$, with $I_{\nu}$ the intensity per unit frequency, is conserved along the geodesic (Lindquist 1966), so that the source specific intensity $I_{\nu}^{\mathrm{em}}$ is related to the observed one $I_{\nu}^{\mathrm{obs}}$:
@@
Even though plasma-physics phenomena, such as magnetic reconnection, can accelerate a population of electrons in power-law distributions, we assume here, to simplify, that the entire disc emits relativistic thermal synchrotron radiation; which is reasonable to characterise M87* emission at millimetre wavelengths [^18] The assumption is not appropriate for the infra-red flaring of Sgr A*..
Thus, according to equation (B3) in Vincent et al. (2022), and making the same approximations detailed therein in Appendix A, the relativistic electrons follow the Maxwell-Jüttner distribution at temperature $T_{\mathrm{e}}$, whose emissivity in cgs units is:
@@
$\theta_{\mathrm{B}}$ $=\arccos(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})\>.$ (18.1)
where $n_{\mathrm{e}}$ is the number density of the electrons of mass $m_{\mathrm{e}}$, $e$ the charge of the electrons, $B$ the amplitude of the magnetic field, $\theta_{\mathrm{B}}$ the pitch angle between the unit vector along the direction of emission $\boldsymbol{\bar{K}}$ and the unit magnetic field vector $\boldsymbol{\bar{B}}$, both measured in the comoving frame of the emitting plasma [^19] See Appendix C.3 for a rigorous definition of $\boldsymbol{\bar{K}}$ and $\boldsymbol{\bar{B}}$., and $\Theta_{\mathrm{e}}\coloneqq k_{\mathrm{B}}T_{\mathrm{e}}/m_{\mathrm{e}}c^{2}$, $k_{B}$ being the Boltzmann constant.
Because the disc of our simulation is geometrically and optically thin, the radiative transfer equation reduces to:
@@
In order to study the effects of the emission process on the simulated images, we assume that the physical quantities $n_{\mathrm{e}}$, $\Theta_{\mathrm{e}}$ and $B$ follow power-law radial distributions with indices $\alpha$, $\beta$ and $\gamma$ varying in ranges including the values obtained by 3D GRMHD simulations (Cho et al. 2023; Chatterjee et al. 2023) and in fair agreement with the ones set by Desire et al. (2025):
[…]
Substituting these power-laws in the emissivity, introducing the indices $i_{1}=\alpha-2\beta$ and $i_{2}=\gamma+2\beta$, gives:
$j_{\nu}$ $\approx\eta\frac{\nu_{\mathrm{em}}\text{\scriptsize[GHz]}}{230}\left(\frac{r}{r_{\mathrm{inner}}}\right)^{-i_{1}}\exp\left[-\zeta\sqrt[3]{\frac{\nu_{\mathrm{em}}\text{\scriptsize[GHz]}}{230\sin\theta_{\mathrm{B}}}}\left(\frac{r}{r_{\mathrm{inner}}}\right)^{i_{2}/3}\right],$ (21)
@@
The value of $\zeta$, justified in detail in Appendix B of Vincent et al. (2022), is chosen in order to match the predicted characteristics of the accretion flow around M87*. The normalisation $\eta>0$, ensures that the 230 GHz radiative flux density is of the order of 0.5 Jy for a mean profile ($\zeta$=3, $\alpha$=1, $\beta$=2) as suggested by the EHT data (The Event Horizon Telescope Collaboration 2019b; Wielgus et al. 2020; The Event Horizon Telescope Collaboration 2024). Hence, our simulations only result in a correct flux density for this particular set of parameters. Nevertheless, since the normalization $\eta$ does not influence the conclusions of this work, we keep its value constant for simplicity.
Table 2: Parameters of the accretion disc
@@ 5 Adaptive backward ray-tracing
To simulate our images we use the ray-tracing code GYOTO (General RelativitY Orbit Tracer of the Observatoire de Paris, Vincent et al. 2011), collecting the specific intensity $I_{\nu}$ along null geodesics which are shot from pixels making up the screen of a distant observer towards the compact object. The resulting image is then a map of $I_{\nu}$ on the observer sky.
@@
Keeping track of the equatorial crossings, it is possible to decompose the simulated image into different $n$ orders, properly defined for light rays in Section 2: the $n=0$ image is the map of the specific intensity obtained along the part of geodesic integrated before the second equatorial plane encounter, while the $n=1$ image is the map of $I_{\nu}$ accumulated between the second and the third crossing. We restrict our analyses to the zeroth and first orders so that the geodesic integration leading to the $n\geq 2$ images is not performed.
@@
We define a polar coordinate system on the observer’s screen, that is, on the 2D simulated image, so that every point on it is described by its distance $b$ from the centre $O$ and by the positive angle $\psi$ from the horizontal axis passing from $O$, see Figure 4.
@@
For each metric of Section 3, we produce adaptively ray-traced images based on the identification of the projection of the equatorial horizon and lensing bands introduced in Section 2, adapting the resolution for different image orders which are characterised by different scales. As explained in Gelles et al. (2021), this is an efficient way to explore the properties of computationally-expensive photon rings. To do that, we first simulate the image of a disc with constant emission and without spectral redshift effects ($\nu_{\mathrm{em}}=\nu_{\mathrm{obs}}$, $\zeta=0$ and $i_{1}=0$) on pixels separated by 0.1 µas along 14 equally spaced polar angular directions on sky. The disc is taken to have a very large effective outer radius in this case. Then, since the order of a null geodesic is related to the number of equatorial crossings, the structure of the image can be distinguished thanks to the number of impacts of the photon with the disc, itself linked to the value of $I_{\nu}$. In other words, for each angular direction, the boundary of the projected horizon is given by the further pixels from the centre having a null intensity, as no emission originates from the interior of the black hole by definition, and the edges of the $n=1$ lensing band are defined by the closest and furthest pixels cumulating the maximum intensity, since their geodesic collect twice the constant emission of the disc. Afterwards, the 14 computed values on the edges are connected with an interpolated curve. Next, for the actual emission profiles of Section 4, the $n=0$ image, computed for the pixels in the mask extending outside the apparent horizon until $r_{\mathrm{outer}}$ and the $n=1$ image, computed only inside the mask corresponding to the $n=1$ lensing band, with margins of 0.1 µas around the desired zone in order to be sure not to lose some needed pixels because of a bad initial resolution, are traced with adaptively chosen resolutions $\mathrm{res}^{b,\>n=0,1}$.
@@
Our images are simulated at the two [^21] At 86 GHz absorption effects cannot be neglected. highest observing frequencies of the ngEHT and BHEX, 230 and 345 GHz (Doeleman et al. 2023; Johnson et al. 2024), and the screen orientation is chosen in such a way that the direction of the projected angular-momentum vector of the disc is aligned with the vertical axis. The field of view and the integration variables are chosen in order to have a reasonable compromise between a good image resolution and a modest computing time. The parameters of our simulations are given in Tables 1, 2 and 3.
Table 3: Parameters of the simulation
@@
Figure 5: Dependence on the metric parameters $\epsilon$, $a_{1}$ of the event horizon, the photon sphere, the critical curve and the innermost stable circular orbit. The curves depending on $\epsilon$ or $a_{1}$ are traced at $a_{1}=0$ or $\epsilon=0$ respectively. Green lines indicate corresponding radii in a Schwarzschild spacetime.

### v2

@@ 4 Accretion disc model
In this paper, the compact object is surrounded by an infinitely geometrically thin accretion disc orbiting in the equatorial plane taken to be at $\theta=\pi/2$ without any loss of generality. Our disc extends from an inner radius, $r_{\mathrm{inner}}$, given by the radial position of the event horizon, to a sufficiently large outer radius, $r_{\mathrm{outer}}$, whose contribution to the image formation is negligible for our choice of quickly decreasing emission profiles.
@@ 4.1 Motion of the emitting material
We supposed that the emitters of mass $m_{\mathrm{e}}>0$ can have three possible types of dynamics: the particles in the disc rotate around the black hole with a Keplerian velocity, they fall radially towards the central singularity, or they possess a more realistic mixed velocity. A sketch of the two extreme velocity fields is reproduced in Figure 3. We defined the specific energy, $\varepsilon$, and specific angular momentum, $\ell$, as
@@ 4.1.1 Keplerian velocity
If the disc is Keplerian, the accreted matter follows a circular geodesic motion above the innermost stable circular orbit (ISCO), below which we assumed that it spirals down along time-like geodesics with constants of motion, $\varepsilon$ and $\ell$, equal to those of the ISCO (Cunningham 1975). A general time-like 4-velocity, $u^{\,\mu}=p^{\,\mu}/m_{\mathrm{e}}=(u^{t},u^{r},0,u^{\varphi})$, can be expressed with the conserved quantities and via the normalisation $u^{\,\mu}u_{\mu}=-1$. Noting by primes the derivatives with respect to $r$, the expressions of the specific energy and z-component of the angular momentum for the RZ metric are then (Cárdenas-Avendaño et al. 2019)
@@
The last stable orbit being marginally stable, the second derivative of the effective potential, $V^{\prime\prime}_{\mathrm{eff}}$, must vanish at the radius of the ISCO, $r_{\mathrm{ms}}$, which is then the real root of the following equation with no analytical solution (Rezzolla & Zhidenko 2014):
@@ 4.1.2 Radial infall
When the accreted matter is in radial free-fall towards the central compact object, the particle evolves in a constant-$\theta$ plane with vanishing specific angular momentum, $\ell$. The conservation of the specific energy, $\varepsilon$, and the normalisation of the 4-velocity lead to (Kocherlakota & Rezzolla 2022):
@@
We also neglected the radial velocity at infinity so that
@@ 4.1.3 Mixed velocity
Realistic RIAF discs are neither totally radially infalling nor Keplerian, and according to more sophisticated GRMHD simulations (Narayan et al. 2012), they likely consist of a mixed velocity comprising a radial component and a sub-Keplerian part (Yuan et al. 2022; Begelman et al. 2022). Moreover, the sub-Keplerian motion is expected to be predominant for the disc surrounding M87* (Chatterjee et al. 2023).
To define the sub-Keplerian velocity, $u_{\mathrm{sub}}$, we followed the same reasoning of Paragraph 4.1.1, but we multiplied the quantity $\ell/\varepsilon$ by a ‘sub-Keplerianity’ factor, $\xi\in(0,1]$. Thus, above the ISCO
$\varepsilon_{\mathrm{sub}}=\sqrt{\frac{N^{3}(r)}{N(r)-\xi^{2}rN^{\prime}(r)}}\>,\quad\ell_{\mathrm{sub}}=\xi\sqrt{\frac{r^{3}N^{\prime}(r)}{N(r)-\xi^{2}rN^{\prime}(r)}}\>,$
whereas below the ISCO $\varepsilon_{\mathrm{sub}}=\varepsilon_{\mathrm{sub}}(r_{\mathrm{ms}})$ and $\ell_{\mathrm{sub}}=\ell_{\mathrm{sub}}(r_{\mathrm{ms}})$. Then, introducing the angular velocity, $\Omega=u^{\varphi}/u^{t}$, [^17] In spherical symmetric space-times, $\Omega_{\mathrm{rad}}$ is always zero. and the parameters $0\leq\omega_{r}\leq 1$ and $0\leq\omega_{\varphi}\leq 1$, we linearly combined the sub-Keplerian and radial 4-velocities (Pu et al. 2016; Cárdenas-Avendaño et al. 2023):
@@
Finally, $u^{t}_{\mathrm{mix}}$ was found by unit-normalisation and $u^{\varphi}_{\mathrm{mix}}=\Omega_{\mathrm{mix}}u^{t}_{\mathrm{mix}}$.
We note that this realistic mixed velocity does not correspond to a geodesic motion contrary to the purely Keplerian and radial ones, which are retrieved, respectively, by imposing $\omega_{r}=\omega_{\varphi}=1$ with $\xi=1$ and $\omega_{r}=\omega_{\varphi}=0$. In other words, the mixed velocity prescription can account for non-gravitational forces, which are expected to affect the emitters’ trajectory.
@@ 4.2 Synchrotron radiation
We assumed the disc to be optically and geometrically thin, which means considering the optical depth to be idealistically null. The first hypothesis is only an approximation for the present EHT images of M87* and particularly for the photon rings observations, for which the opacity effects may play a role (Vincent et al. 2022). The assumption holds much better at 345 GHz, as absorption effects decrease with increasing frequency, and so it is reasonable for the observations at higher frequency of the ngEHT (Doeleman et al. 2023) or BHEX (Johnson et al. 2024). In the absence of absorption, the quantity $I_{\nu}/\nu^{3}$, with $I_{\nu}$ the intensity per unit frequency, is conserved along the geodesic (Lindquist 1966) so that the source specific intensity, $I_{\nu}^{\mathrm{em}}$, is related to the observed intensity, $I_{\nu}^{\mathrm{obs}}$:
@@
Even though plasma-physics phenomena, such as magnetic reconnection, can accelerate a population of electrons in power-law distributions, we assumed, to simplify, that the entire disc emits relativistic thermal synchrotron radiation; which is reasonable to characterise M87* emission at millimetre wavelengths. [^18] The assumption is not appropriate for the infra-red flaring of Sgr A*.
Thus, according to equation (B3) in Vincent et al. (2022), and making the same approximations detailed therein in Appendix A, the relativistic electrons follow the Maxwell-Jüttner distribution at temperature $T_{\mathrm{e}}$, whose emissivity in cgs units is
@@
$\theta_{\mathrm{B}}$ $=\arccos(\boldsymbol{\bar{K}}\cdot\boldsymbol{\bar{B}})\>,$ (18.1)
where $n_{\mathrm{e}}$ is the number density of the electrons of mass $m_{\mathrm{e}}$, $e$ the charge of the electrons, $B$ the amplitude of the magnetic field, $\theta_{\mathrm{B}}$ the pitch angle between the unit vector along the direction of emission, $\boldsymbol{\bar{K}}$, and the unit magnetic field vector, $\boldsymbol{\bar{B}}$, both measured in the comoving frame of the emitting plasma, [^19] A rigorous definition of $\boldsymbol{\bar{K}}$ and $\boldsymbol{\bar{B}}$ is given in Appendix C.3. and $\Theta_{\mathrm{e}}\coloneqq k_{\mathrm{B}}T_{\mathrm{e}}/m_{\mathrm{e}}c^{2}$, $k_{B}$ being the Boltzmann constant.
Because the disc of our simulation is geometrically and optically thin, the radiative transfer equation reduces to
@@
In order to study the effects of the emission process on the simulated images, we assumed that the physical quantities, $n_{\mathrm{e}}$, $\Theta_{\mathrm{e}}$, and $B$, follow power-law radial distributions with indices $\alpha$, $\beta$, and $\gamma$ varying in ranges including the values obtained by 3D GRMHD simulations (Cho et al. 2023; Chatterjee et al. 2023) and in fair agreement with the ones set by Desire et al. (2025):
$\begin{aligned} n_{e}&\propto r^{-\alpha},\quad\alpha\in[1;1.5]\\ \Theta_{e}&\propto r^{-\beta},\quad\beta\,\in[1;2]\\ B&\propto r^{-\gamma},\quad\gamma=1\>.\end{aligned}$ (20)
Substituting these power-laws in the emissivity, introducing the indices $i_{1}=\alpha-2\beta$ and $i_{2}=\gamma+2\beta$, gives
$j_{\nu}$ $\approx\eta\frac{\nu_{\mathrm{em}}\text{\scriptsize[GHz]}}{230}\left(\frac{r}{r_{\mathrm{inner}}}\right)^{-i_{1}}\hskip-2.84544pt\exp\left[-\zeta\sqrt[3]{\frac{\nu_{\mathrm{em}}\text{\scriptsize[GHz]}}{230\sin\theta_{\mathrm{B}}}}\left(\frac{r}{r_{\mathrm{inner}}}\right)^{i_{2}/3}\right],$ (21)
@@
The value of $\zeta$, justified in detail in Appendix B of Vincent et al. (2022), was chosen in order to match the predicted characteristics of the accretion flow around M87*. The normalisation, $\eta>0$, ensures that the 230 GHz radiative flux density is of the order of 0.5 Jy for a mean profile ($\zeta$=3, $\alpha$=1, $\beta$=2) as suggested by the EHT data (The Event Horizon Telescope Collaboration 2019b; Wielgus et al. 2020; The Event Horizon Telescope Collaboration 2024). Hence, our simulations only result in a correct flux density for this particular set of parameters. Nevertheless, since the normalisation does not influence the conclusions of this work, we kept the value of $\eta$ constant for simplicity.
Table 2: Parameters of the accretion disc.
@@ 5 Adaptive backward ray-tracing
To simulate our images we used the ray-tracing code GYOTO (General RelativitY Orbit Tracer of the Observatoire de Paris, Vincent et al. 2011), collecting the specific intensity, $I_{\nu}$, along null geodesics that are shot from pixels making up the screen of a distant observer towards the compact object. The resulting image is then a map of $I_{\nu}$ on the observer sky.
@@
Keeping track of the equatorial crossings, it is possible to decompose the simulated image into different $n$ orders, properly defined for light rays in Section 2: the $n=0$ image is the map of the specific intensity obtained along the part of geodesic integrated before the second equatorial plane encounter, while the $n=1$ image is the map of $I_{\nu}$ accumulated between the second and the third crossing. We restricted our analyses to the zeroth and first orders so that the geodesic integration leading to the $n\geq 2$ images was not performed.
@@
We defined a polar coordinate system on the observer’s screen, that is, on the 2D simulated image, so that every point on it is described by its distance, $b$, from the centre, $O$, and by the positive angle, $\psi$, from the horizontal axis passing from $O$ (see Figure 4).
@@
For each metric of Section 3, we produced adaptively ray-traced images based on the identification of the projection of the equatorial horizon and lensing bands introduced in Section 2, adapting the resolution for different image orders which are characterised by different scales. As explained in Gelles et al. (2021), this is an efficient way to explore the properties of computationally expensive photon rings. To do that, we first simulated the image of a disc with constant emission and without spectral redshift effects ($\nu_{\mathrm{em}}=\nu_{\mathrm{obs}}$, $\zeta=0$ and $i_{1}=0$) on pixels separated by 0.1 µas along 14 equally spaced polar angular directions on sky. The disc was taken to have a very large effective outer radius in this case. Then, since the order of a null geodesic is related to the number of equatorial crossings, the structure of the image could be distinguished thanks to the number of impacts of the photon with the disc, itself linked to the value of $I_{\nu}$. In other words, for each angular direction, the boundary of the projected horizon was given by the further pixels from the centre having a null intensity, as no emission originates from the interior of the black hole by definition, and the edges of the $n=1$ lensing band were defined by the closest and furthest pixels cumulating the maximum intensity, since their geodesic collected twice the constant emission of the disc. Afterwards, the 14 computed values on the edges were connected with an interpolated curve. Next, for the actual emission profiles of Section 4, the $n=0$ image, computed for the pixels in the mask extending outside the apparent horizon until $r_{\mathrm{outer}}$ and the $n=1$ image, computed only inside the mask corresponding to the $n=1$ lensing band, with margins of 0.1 µas around the desired zone in order to be sure not to lose some needed pixels because of a bad initial resolution, were traced with adaptively chosen resolutions, $\mathrm{res}^{b,\>n=0,1}$.
@@
Our images were simulated at the two [^21] At 86 GHz absorption effects cannot be neglected. highest observing frequencies of the ngEHT and BHEX, 230 and 345 GHz (Doeleman et al. 2023; Johnson et al. 2024), and the screen orientation was chosen in such a way that the direction of the projected angular-momentum vector of the disc is aligned with the vertical axis. The field of view and the integration variables were chosen in order to have a reasonable compromise between a good image resolution and a modest computing time. The parameters of our simulations are given in Tables 1, 2 and 3.
Table 3: Parameters of the simulation.
@@
Figure 5: Dependence on the metric parameters $\epsilon$, $a_{1}$ of the event horizon, the photon sphere, the critical curve, and the ISCO. The curves depending on $\epsilon$ or $a_{1}$ are traced at $a_{1}=0$ or $\epsilon=0$ respectively. Green lines indicate corresponding radii in a Schwarzschild space-time.